OpenStudy (dinorap1):

Please help! Will give medal! 4. What is the vertex form of the equation below? y=x^2+4x-3

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \[-\frac{4}{2\times 1}=-2\]

OpenStudy (anonymous):

second coordinate of the vertex is is what you get when you replace \(x\) by \(-2\)

OpenStudy (anonymous):

once you have your vertex \((h,k)\) vertex form is \[y=(x-h)^2+k\]

OpenStudy (dinorap1):

@satellite73 Oh, I get it. Soo... you replace x with -2, and you get y=-7, right?

OpenStudy (dinorap1):

So does it become y=(x-2)^2-7?

OpenStudy (dinorap1):

I'm assuming that's the answer... thanks for the help @satellite73 ! Sorry for not being here when you replied! :(